OpenStudy (anonymous):

Find a function that models the area A of a square in terms of the diagonal d.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

From Geometry, recall that the area of a rhombus can be found by the formula: Area = (1/2)*d1*d2, where d1 and d2 are the lengths of the two diagonals.

Directrix (directrix):

A square is a type of rhombus so that formula applies to the square.

Directrix (directrix):

Also, recall from Geometry that the diagonals of a square are congruent. So, in the area formula: Area = (1/2)*d1*d2, where d1 and d2, d1 = d2.

Directrix (directrix):

Let the diagonal of the square in question just be d. Area = (1/2) d * d = ? @ninab731

OpenStudy (anonymous):

So then would then our answer be just 1/2d^2

Directrix (directrix):

Yes. Write that as an equation: A = (1/2) d^2.

OpenStudy (anonymous):

Cool thanks :) so now about function that models the diagonal d of a square in terms of the perimeter

OpenStudy (anonymous):

I know that perimeter of a square is all the sides

OpenStudy (anonymous):

@Directrix one more ? Please :)

Directrix (directrix):

From Geometry, recall that the diagonal of a square has measure √2 * a side of a square. This is a result of applying the Pythagorean theorem. So, d = √2 * a side of a square